OpenStudy (anonymous):

b+3 over e equals f over 2 solve for e

OpenStudy (bahrom7893):

not really: (b+3)/e = f/2 Multiply everything by 2e: 2(b+3) = ef Divide everything by f: e = (2(b+3))/f